Globus Autos sells a single product. 8,400 units were sold resulting in $82,000 of sales revenue, $22,000 of variable costs, and $12,000 of fixed costs. If variable costs decrease by $1.2 per unit, the new margin of safety is ________.
Answer #1

New margin of safety = 67,958.90 or 67959 in percentage 82.88% or 83%

Working

A Sales price [82000/8400] $               9.76
B Variable cost per unit {[22000/8400]-1.2} $               1.42
C=A-B Contribution per unit $               8.34
D Fixed cost $    12,000.00
E=D/C Breakeven Sales in units                 1,438
F= E x A Breakeven Sales in Dollars $    14,041.10
G Actual sales $    82,000.00
H=G-F Margin of Safety in Dollars $    67,958.90
I=H/G Margin of safety percentage 82.88%